# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 2.4. Polynomial chaos expansion
(2.12) $\begin{split}Z_{N,\beta_{N}}^{\omega}=&{\mathbf{E}}\Big{[}\prod\limits_{n=1}^{N}\Big{(}1+\zeta_{n}\mathbbm{1}_{\{S_{n}=0\}}\Big{)}\Big{]}=1+\sum\limits_{k=1}^{N}\beta_{N}^{k}\sum\limits_{1\leq n_{1}<\cdots<n_{k}\leq N}\prod\limits_{i=1}^{k}p_{n_{i-1},n_{i}}(0)\zeta_{n_{i}},\end{split}$
@@
(2.13) $p_{m,n}(x,y):={\mathbf{P}}(S_{n}=y|S_{m}=x),\quad\text{for}~{}m<n,$
@@
(2.14) $\begin{split}&\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi):=\iint_{{\mathbbm{R}}^{2}}\varphi(x)\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(x,y)\psi(y)\\ =&\sqrt{N}\sum\limits_{u,v\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}}\int_{u}^{u+\frac{1}{\sqrt{N}}}\varphi(x){\mathrm{d}}x\Big{(}\widetilde{Z}_{0,N}^{\omega,\beta_{N}}\big{(}u\sqrt{N},v\sqrt{N}\big{)}\Big{)}\int_{v}^{v+\frac{1}{\sqrt{N}}}\psi(y){\mathrm{d}}y\\ =&q_{0,N}^{N}(\varphi,\psi)+\frac{1}{\sqrt{N}}\sum\limits_{r=1}^{\infty}\sum_{0<n_{1}<\cdots<n_{r}<N}q_{0,n_{1}}^{N}(\varphi,0)\zeta_{n_{1}}\Big{(}\prod\limits_{j=2}^{r}p_{n_{j-1},n_{j}}(0)\zeta_{n_{j}}\Big{)}q_{n_{r},N}^{N}(0,\psi),\end{split}$
@@
(2.15) $q_{0,n}^{N}(\varphi,x)$ $:=\sum\limits_{u\in{\mathbbm{Z}}}\varphi_{N}(u)p_{n}(x-u),$
(2.16) $q_{m,N}^{N}(y,\psi)$ $:=\sum\limits_{v\in{\mathbbm{Z}}}p_{m,N}(v-y)\psi_{N}(v),$
(2.17) $q_{0,N}^{N}(\varphi,\psi)$ $:=\frac{1}{\sqrt{N}}\sum\limits_{u,v\in{\mathbbm{Z}}}\varphi_{N}(u)p_{N}(v-u)\psi_{N}(v),$
@@
(2.18) $\varphi_{N}(x):=\int_{x}^{x+1}\varphi\Big{(}\frac{t}{\sqrt{N}}\Big{)}{\mathrm{d}}t,\quad\psi_{N}(y):=\int_{y}^{y+1}\psi\Big{(}\frac{s}{\sqrt{N}}\Big{)}{\mathrm{d}}s.$
@@
(2.19) $\begin{split}{\mathbbm{E}}\left[(\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi))^{2}\right]&=\left(q_{0,N}^{N}(\varphi,\psi)\right)^{2}+\\ \frac{1}{N}&\sum\limits_{r=1}^{\infty}(\sigma_{N}^{2})^{r}\sum\limits_{0<n_{1}<\cdots<n_{r}<N}\left(q_{0,n_{1}}^{N}(\varphi,0)\right)^{2}\Big{(}\prod\limits_{j=2}^{r}u(n_{j}-n_{j-1})\Big{)}\left(q_{n_{r},N}^{N}(0,\psi)\right)^{2},\end{split}$
@@
(2.20) $u(n):=p_{n}^{2}(0)\sim\frac{1}{2\pi n}.$
A powerful tool for computing the second moment (2.19) is the so-called Dickman subordinator, which we now introduce in the following subsection.
2.5. Dickman subordinator and second moment computations
We first introduce a renewal process $\iota^{(N)}=\{\iota^{(N)}_{1},\iota^{(N)}_{2},\cdots\}$ on ${\mathbbm{N}}\cup\{0\}$ with one-step distribution
(2.21) ${\mathbf{P}}\big{(}\iota^{(N)}_{1}=n\big{)}=\frac{u(n)}{R_{N}}\mathbbm{1}_{\{1,\cdots,N\}}(n),$
where $R_{N}$ is the expected overlap (see (1.11)).
In view of (2.19) above, if the first and last return times to 00 for the random walk are denoted by $m$ and $n$ respectively with $m<n$, then we have that by the renewal structures of the random walk and (2.21),
(2.22) $\begin{split}&(\sigma_{N}^{2})^{2}u(n-m)+\sum\limits_{k=1}^{\infty}(\sigma_{N}^{2})^{k+2}\sum\limits_{n_{0}:=m<n_{1}<\cdots<n_{k}<n=:n_{k+1}}\prod\limits_{j=1}^{k+1}u(n_{j}-u_{j-1})\\ =&\sigma_{N}^{2}\sum\limits_{k=1}^{\infty}(\lambda_{N})^{k}{\mathbf{P}}(\iota^{(N)}_{k}=n-m),\quad\text{where}~{}\lambda_{N}:=\sigma_{N}^{2}R_{N}=1+\frac{\vartheta+o(1)}{\log N}.\end{split}$
Hence, to analyze (2.19), in particular, (2.22), it turns out that we can investigate the renewal process $\iota^{(N)}$, which has been intensively studied in [CSZ19a]. We cite the following result.
Proposition 2.1 ([CSZ23], Lemma 3.3).
@@ 2.5. Dickman subordinator and second moment computations
where $Y_{s}$ is a pure jump Lévy process with Lévy measure $\frac{1}{t}\mathbbm{1}_{(0,1)}(t){\mathrm{d}}t$ called the Dickman subordinator, and the convergence in distribution holds on the space of càdlàg paths equipped with the Skorohod topology. Moreover, the density of $Y_{s}$ is known (cf. [CSZ19a, Theorem 1.1]), which is
[…]
for all $s\in(0,+\infty)$, where Γ()normal-$\Gamma(\cdot)$ is the Gamma function and $\gamma=-\int_{0}^{\infty}\log ue^{-u}{\mathrm{d}}u$ is the Euler-Mascheroni constant.
To include the case that there is only one return to 00 (i.e., $m=n$ in (2.22)), we introduce the following quantity, which plays a central role throughout the paper. Let
(2.24) $\overline{U}_{N}(n):=\begin{cases}\sigma_{N}^{2},&\quad\text{if}~{}n=0,\\ \displaystyle(\sigma_{N}^{2})^{2}u(n)+\sum\limits_{k=1}^{\infty}(\sigma_{N}^{2})^{k+2}\sum\limits_{n_{0}:=0<n_{1}<\cdots<n_{k}<n=:n_{k+1}}\prod\limits_{j=1}^{k+1}u(n_{j}-n_{j-1}),&\quad\text{if}~{}n\geq 1.\end{cases}$
We then collect some useful properties of $\overline{U}_{N}(n)$ from [CSZ19a, Section 1-3]. Given $\vartheta$ in (1.20), let
(2.25) $G_{\vartheta}(t):=\int_{0}^{\infty}e^{\vartheta s}\alpha_{s}(t){\mathrm{d}}s,$
where $\alpha_{s}(\cdot)$ is the density of the Dickman subordinator $Y_{s}$ (see Proposition 2.1).
  • By [CSZ19a, Theorem 1.4], given $T>0$, there exists some constant $C>0$, such that
    (2.26) $\overline{U}(n)\leq\frac{C}{N}G_{\vartheta}\left(\frac{n}{N}\right),\quad\text{uniformly for all}~{}1\leq n\leq TN.$
    Furthermore, if $\delta\in(0,1)$ is also given, then
    (2.27) $\overline{U}_{N}(n)=\frac{2\pi}{N}\left(G_{\vartheta}\left(\frac{n}{N}\right)+o_{N}(1)\right),\quad\text{uniformly for all}~{}\delta N\leq n\leq TN.$
  • By [CSZ19a, Proposition 1.6], the asymptotics of $G_{\vartheta}(t)$ is known, that is,
    (2.28) $G_{\vartheta}(t)\sim\frac{1}{t}\Big{(}\log\frac{1}{t}\Big{)}^{-2}\quad\text{and}\quad\int_{0}^{t}G_{\vartheta}(s){\mathrm{d}}s\sim\Big{(}\log\frac{1}{t}\Big{)}^{-1},\quad\text{as}\quad t\to 0.$
@@
It is not surprising that the Dickman subordinator and the function $\overline{U}(\cdot)$ appear both in [CSZ23] and our paper. If we perform a polynomial chaos expansion to $\overline{\mathcal{Z}}_{N}$ in Subsection 2.2 and compute its second moment, then we have
$\begin{split}{\mathbbm{E}}\left[(\overline{\mathcal{Z}}_{N})^{2}\right]&=\left(q_{0,N}^{N}(\varphi,\psi)\right)^{2}+\\ \frac{1}{N}&\sum\limits_{r=1}^{\infty}(\sigma_{N}^{2})^{r}\sum\limits_{0<n_{1}<\cdots<n_{r}<N\atop x_{1},\cdots,x_{r}\in{\mathbbm{Z}}^{2}}\left(q_{0,n_{1}}^{N}(\varphi,x_{1})\right)^{2}\Big{(}\prod\limits_{j=2}^{r}p_{n_{j-1},n_{j}}(x_{j-1},x_{j})^{2}\Big{)}\left(q_{n_{r},N}^{N}(x_{r},\psi)\right)^{2},\end{split}$
@@
Next, we introduce a pair of a renewal process and a random walk $(\overline{\iota}^{(N)}_{k},\overline{S}^{(N)}_{k})_{k\geq 0}$ on ({$({\mathbbm{N}}\cup\{0\})\times{\mathbbm{Z}}^{2}$ starting at $(0,0)$ with one-step probability
@@
(2.29) $g_{t}(\varphi,a)$ $:=\int_{\mathbbm{R}}\varphi(x)g_{t}(a-x){\mathrm{d}}x,$
(2.30) $g_{t}(b,\psi)$ $:=\int_{\mathbbm{R}}g_{t}(y-b)\psi(y){\mathrm{d}}y,$
(2.31) $g_{t}(\varphi,\psi)$ :=×$:=\int_{{\mathbbm{R}}\times{\mathbbm{R}}}\varphi(x)g_{t}(y-x)\varphi(y){\mathrm{d}}x{\mathrm{d}}y.$
@@
(2.32) $\lim\limits_{N\to\infty}{\mathbbm{E}}\Big{[}\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)\Big{]}$ $=g(\varphi,\psi),$
(2.33) $\lim\limits_{N\to\infty}{\mathbbm{E}}\Big{[}\Big{(}\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)\Big{)}^{2}\Big{]}$ $=g(\varphi,\psi)^{2}+{\mathcal{V}}_{1}^{\vartheta}(\varphi,\psi)$
@@
(2.34) ${\mathcal{V}}_{1}^{\vartheta}(\varphi,\psi)=\iiiint_{{\mathbbm{R}}^{\otimes 4}}\varphi(x)\varphi(x^{\prime})K_{1}^{\vartheta}(x,x^{\prime};y,y^{\prime})\psi(y)\psi(y^{\prime}){\mathrm{d}}x{\mathrm{d}}x^{\prime}{\mathrm{d}}y{\mathrm{d}}y^{\prime}\leq 2\pi\|\varphi\|_{\infty}^{2}\|\psi\|^{2}_{\infty}\int_{0}^{1}G_{\vartheta}(u){\mathrm{d}}u$
@@
(2.35) $K_{t}^{\vartheta}(x,x^{\prime};y,y^{\prime})$ $:=2\pi\iint_{0<r<s<t}g_{r}(x)g_{r}(x^{\prime})G_{\vartheta}(s-r)g_{t-s}(y)g_{t-s}(y^{\prime}){\mathrm{d}}r{\mathrm{d}}s.$
@@
The limit (2.32) directly follows by the local limit theorem and a Riemann sum approximation.
For (2.33), we only need to treat the second line of (2.19). We have that
(2.36) $\begin{split}q_{0,n}^{N}(\varphi,0)&=\sum\limits_{u\in{\mathbbm{Z}}}\sqrt{N}\int_{\frac{u}{\sqrt{N}}}^{\frac{u+1}{\sqrt{N}}}\varphi(x){\mathbf{P}}(S_{n}=u){\mathrm{d}}x=\int_{\mathbbm{R}}\varphi(x)g_{\frac{n}{N}}(x){\mathrm{d}}x+o_{N}(1)\\ q_{n,N}^{N}(0,\psi)&=\sum\limits_{v\in{\mathbbm{Z}}}\sqrt{N}\int_{\frac{v}{\sqrt{N}}}^{\frac{v+1}{\sqrt{N}}}\psi(y){\mathbf{P}}(S_{N-n}=v){\mathrm{d}}y=\int_{\mathbbm{R}}\varphi(y)g_{1-\frac{n}{N}}(y){\mathrm{d}}y+o_{N}(1).\end{split}$
@@
By (2.26), the contribution from $n\leq\varepsilon N$ and $n\geq(1-\varepsilon)N$ is bounded by $C_{\varphi,\psi}\int_{[0,\varepsilon]\cup[1-\varepsilon,1]}G_{\vartheta}(t){\mathrm{d}}t$, which converges to 00 as $\varepsilon\to 0$. The proof is completed. ∎
@@
$\sum\limits_{r=1}^{\infty}(\widehat{h})^{r}\idotsint\limits_{0<t_{1}<\cdots<t_{r}<1}g_{t_{1}}(\varphi,0)\Big{(}\prod\limits_{i=2}^{r}\frac{1}{\sqrt{t_{i}-t_{i-1}}}\Big{)}g_{1-t_{r}}(\psi,0){\mathrm{d}}t_{1},\dots\mathrm{d}t_{r},$
which is convergent by [CSZ16, Lemma B.2]. Note that as long as $h_{N}=o((\log N)^{-1})$, the asymptotics of $\sigma_{N}^{2}$ in (1.20) remains unchanged. Since $h_{N}$ only affects the mean, we can set $h_{N}\equiv 0$ without loss of generality.
2.6. Coarse graining procedure

### v3

@@ 2.4 Polynomial chaos expansion
$\begin{split}Z_{N,\beta_{N}}^{\omega}=&{\mathbf{E}}\Big[\prod\limits_{n=1}^{N}\Big(1+\zeta_{n}\mathbbm{1}_{\{S_{n}=0\}}\Big)\Big]=1+\sum\limits_{k=1}^{N}\sum\limits_{1\leq n_{1}<\cdots<n_{k}\leq N}\prod\limits_{i=1}^{k}p_{n_{i-1},n_{i}}(0)\zeta_{n_{i}},\end{split}$ (2.12)
@@
$p_{m,n}(x,y):={\mathbf{P}}(S_{n}=y|S_{m}=x),\quad\text{for}~m<n,$ (2.13)
@@
$\begin{split}&\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi):=\iint_{{\mathbbm{R}}^{2}}\varphi(x)\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(x,y)\psi(y)\\ =&\sqrt{N}\sum\limits_{u,v\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}}\int_{u}^{u+\frac{1}{\sqrt{N}}}\varphi(x){\mathrm{d}}x\Big(\widetilde{Z}_{0,N}^{\omega,\beta_{N}}\big(u\sqrt{N},v\sqrt{N}\big)\Big)\int_{v}^{v+\frac{1}{\sqrt{N}}}\psi(y){\mathrm{d}}y\\ =&q_{0,N}^{N}(\varphi,\psi)+\frac{1}{\sqrt{N}}\sum\limits_{r=1}^{\infty}\sum_{0<n_{1}<\cdots<n_{r}<N}q_{0,n_{1}}^{N}(\varphi,0)\zeta_{n_{1}}\Big(\prod\limits_{j=2}^{r}p_{n_{j-1},n_{j}}(0)\zeta_{n_{j}}\Big)q_{n_{r},N}^{N}(0,\psi),\end{split}$ (2.14)
@@
$q_{0,n}^{N}(\varphi,x)$ $:=\sum\limits_{u\in{\mathbbm{Z}}}\varphi_{N}(u)p_{n}(x-u),$ (2.15)
$q_{m,N}^{N}(y,\psi)$ $:=\sum\limits_{v\in{\mathbbm{Z}}}p_{m,N}(v-y)\psi_{N}(v),$ (2.16)
$q_{0,N}^{N}(\varphi,\psi)$ $:=\frac{1}{\sqrt{N}}\sum\limits_{u,v\in{\mathbbm{Z}}}\varphi_{N}(u)p_{N}(v-u)\psi_{N}(v),$ (2.17)
@@
$\varphi_{N}(x):=\int_{x}^{x+1}\varphi\Big(\frac{t}{\sqrt{N}}\Big){\mathrm{d}}t,\quad\psi_{N}(y):=\int_{y}^{y+1}\psi\Big(\frac{s}{\sqrt{N}}\Big){\mathrm{d}}s.$ (2.18)
@@
$\begin{split}&{\mathbbm{E}}\left[(\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi))^{2}\right]=\left(q_{0,N}^{N}(\varphi,\psi)\right)^{2}+\\ \frac{1}{N}\sum\limits_{r=1}^{\infty}&(\sigma_{N}^{2})^{r}\sum\limits_{0<n_{1}<\cdots<n_{r}<N}\left(q_{0,n_{1}}^{N}(\varphi,0)\right)^{2}\Big(\prod\limits_{j=2}^{r}u(n_{j}-n_{j-1})\Big)\left(q_{n_{r},N}^{N}(0,\psi)\right)^{2},\end{split}$ (2.19)
@@
$u(n):=p_{n}^{2}(0)\sim\frac{1}{2\pi n}.$ (2.20)
A powerful tool for computing the second moment (LABEL:varN) is the so-called Dickman subordinator, which we now introduce in the following subsection.
2.5 Dickman subordinator and second moment computations
We first introduce a renewal process $\iota^{(N)}=\{\iota^{(N)}_{1},\iota^{(N)}_{2}$$\cdots\}$ on ${\mathbbm{N}}\cup\{0\}$ with one-step distribution
${\mathbf{P}}\big(\iota^{(N)}_{1}=n\big)=\frac{u(n)}{R_{N}}\mathbbm{1}_{\{1,\cdots,N\}}(n),$ (2.21)
where $R_{N}$ is the expected overlap in (1.11).
In view of (LABEL:varN) above, if the first and last return times to 0 for the random walk are denoted by $m$ and $n$ respectively with $m<n$, then we have that by the renewal structures of the random walk and (2.21),
$\begin{split}&(\sigma_{N}^{2})^{2}u(n-m)+\sum\limits_{k=1}^{\infty}(\sigma_{N}^{2})^{k+2}\sum\limits_{n_{0}:=m<n_{1}<\cdots<n_{k}<n=:n_{k+1}}\prod\limits_{j=1}^{k+1}u(n_{j}-n_{j-1})\\ =&\sigma_{N}^{2}\sum\limits_{k=1}^{\infty}(\lambda_{N})^{k}{\mathbf{P}}(\iota^{(N)}_{k}=n-m),\quad\text{where}~\lambda_{N}:=\sigma_{N}^{2}R_{N}=1+\frac{\vartheta+o(1)}{\log N}.\end{split}$ (2.22)
Hence, to analyze (LABEL:varN), in particular, (2.22), it turns out that we can investigate the renewal process $\iota^{(N)}$, which has been intensively studied in [13]. We cite the following result.
Proposition 2.1 ([15], Lemma 3.3).
@@ 2.5 Dickman subordinator and second moment computations
where $Y_{s}$ is a pure jump Lévy process with Lévy measure $\frac{1}{t}\mathbbm{1}_{(0,1)}(t){\mathrm{d}}t$ called the Dickman subordinator, and the convergence in distribution holds on the space of càdlàg paths equipped with the Skorohod topology. Moreover, the density of $Y_{s}$ is known (cf. [13, Theorem 1.1]), which is
$f_{s}(t)=\begin{cases}\displaystyle\frac{st^{s-1}e^{-\gamma s}}{\Gamma(s+1)},\quad&\text{for}~t\in(0,1],\\ \displaystyle\frac{st^{s-1}e^{-\gamma s}}{\Gamma(s+1)}-st^{s-1}\int_{0}^{t-1}\frac{f_{s}(a)}{(1+a)^{s}}{\mathrm{d}}s,\quad&\text{for}~t\in(1,+\infty),\end{cases}$ (2.23)
for all $s\in(0,+\infty)$, where $\Gamma(\cdot)$ is the Gamma function and $\gamma=-\int_{0}^{\infty}\log ue^{-u}{\mathrm{d}}u$ is the Euler-Mascheroni constant.
To include the case that there is only one return to 0 (i.e., $m=n$ in (2.22)), we introduce the following quantity, which plays a central role throughout the paper. Let
$\overline{U}_{N}(n):=\begin{cases}\sigma_{N}^{2},&~~\text{if}~n=0,\\ \displaystyle(\sigma_{N}^{2})^{2}u(n)+\sum\limits_{k=1}^{\infty}(\sigma_{N}^{2})^{k+2}\sum\limits_{0<n_{1}<\cdots<n_{k}<n}\prod\limits_{j=1}^{k+1}u(n_{j}-n_{j-1}),&~~\text{if}~n\geq 1,\end{cases}$ (2.24)
where we use the convention $n_{0}:=0$ and $n_{k+1}:=n$. We then collect some useful properties of $\overline{U}_{N}(n)$ from [13, Section 1-3].
Given $\vartheta$ in (1.20), let
$G_{\vartheta}(t):=\int_{0}^{\infty}e^{\vartheta s}f_{s}(t){\mathrm{d}}s,$ (2.25)
where $f_{s}(\cdot)$ is the density of the Dickman subordinator $Y_{s}$ as in Proposition 2.1.
  • By [13, Theorem 1.4], given $T>0$, there exists some constant $C>0$, such that
    $\overline{U}_{N}(n)\leq\frac{C}{N}G_{\vartheta}\left(\frac{n}{N}\right),\quad\text{uniformly for all}~1\leq n\leq TN.$ (2.26)
    Furthermore, if $\theta\in(0,1)$ is also given, then
    $\overline{U}_{N}(n)=\frac{2\pi}{N}\left(G_{\vartheta}\left(\frac{n}{N}\right)+o_{N}(1)\right),\quad\text{uniformly for all}~\theta N\leq n\leq TN.$ (2.27)
  • By [13, Proposition 1.6], the asymptotics of $G_{\vartheta}(t)$ is known, that is,
    $G_{\vartheta}(t)\sim\frac{1}{t}\Big(\log\frac{1}{t}\Big)^{-2}\quad\text{and}\quad\int_{0}^{t}G_{\vartheta}(s){\mathrm{d}}s\sim\Big(\log\frac{1}{t}\Big)^{-1},\quad\text{as}\quad t\to 0.$ (2.28)
@@
It is not surprising that the Dickman subordinator and the function $\overline{U}(\cdot)$ appear both in [15] and our paper. If we perform a polynomial chaos expansion to $\overline{{\mathcal{Z}}}_{N}$ in Subsection 2.2 and compute its second moment, then we have
$\begin{split}{\mathbbm{E}}\left[(\overline{{\mathcal{Z}}}_{N})^{2}\right]&=\left(q_{0,N}^{N}(\varphi,\psi)\right)^{2}+\\ \frac{1}{N}\sum\limits_{r=1}^{\infty}&(\sigma_{N}^{2})^{r}\sum\limits_{0<n_{1}<\cdots<n_{r}<N\atop x_{1},\cdots,x_{r}\in{\mathbbm{Z}}^{2}}\left(q_{0,n_{1}}^{N}(\varphi,x_{1})\right)^{2}\Big(\prod\limits_{j=2}^{r}p_{n_{j-1},n_{j}}(x_{j-1},x_{j})^{2}\Big)\left(q_{n_{r},N}^{N}(x_{r},\psi)\right)^{2},\end{split}$
@@
Next, we introduce a pair of a renewal process and a random walk $(\overline{\iota}^{(N)}_{k},\overline{S}^{(N)}_{k})_{k\geq 0}$ on $({\mathbbm{N}}\cup\{0\})\times{\mathbbm{Z}}^{2}$ starting at $(0,0)$ with one-step transition probability
@@
$g_{t}(\varphi,a)$ $:=\int_{\mathbbm{R}}\varphi(x)g_{t}(a-x){\mathrm{d}}x,$ (2.29)
$g_{t}(b,\psi)$ $:=\int_{\mathbbm{R}}g_{t}(y-b)\psi(y){\mathrm{d}}y,$ (2.30)
$g_{t}(\varphi,\psi)$ $:=\int_{{\mathbbm{R}}\times{\mathbbm{R}}}\varphi(x)g_{t}(y-x)\varphi(y){\mathrm{d}}x{\mathrm{d}}y.$ (2.31)
@@
$\lim\limits_{N\to\infty}{\mathbbm{E}}\Big[\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)\Big]$ $=g(\varphi,\psi),$ (2.32)
$\lim\limits_{N\to\infty}{\mathbbm{E}}\Big[\Big(\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)\Big)^{2}\Big]$ $=g(\varphi,\psi)^{2}+{\mathcal{V}}_{1}^{\vartheta}(\varphi,\psi)$ (2.33)
@@
$\begin{split}{\mathcal{V}}_{1}^{\vartheta}(\varphi,\psi)&=\iiiint_{{\mathbbm{R}}^{\otimes 4}}\varphi(x)\varphi(x^{\prime})K_{1}^{\vartheta}(x,x^{\prime};y,y^{\prime})\psi(y)\psi(y^{\prime}){\mathrm{d}}x{\mathrm{d}}x^{\prime}{\mathrm{d}}y{\mathrm{d}}y^{\prime}\\ &\leq 2\pi\|\varphi\|_{\infty}^{2}\|\psi\|^{2}_{\infty}\int_{0}^{1}G_{\vartheta}(u){\mathrm{d}}u\end{split}$ (2.34)
@@
$K_{t}^{\vartheta}(x,x^{\prime};y,y^{\prime})$ $:=2\pi\iint_{0<r<s<t}g_{r}(x)g_{r}(x^{\prime})G_{\vartheta}(s-r)g_{t-s}(y)g_{t-s}(y^{\prime}){\mathrm{d}}r{\mathrm{d}}s.$ (2.35)
@@
The limit (2.32) directly follows by the local limit theorem and a Riemann sum approximation (recall (2.14) and ${\mathbbm{E}}[\zeta_{n}]=0$).
For (2.33), we only need to treat the second line of (LABEL:varN). We have that
$\begin{split}q_{0,n}^{N}(\varphi,0)&=\sum\limits_{u\in{\mathbbm{Z}}}\sqrt{N}\int_{\frac{u}{\sqrt{N}}}^{\frac{u+1}{\sqrt{N}}}\varphi(x){\mathbf{P}}(S_{n}=u){\mathrm{d}}x=\int_{\mathbbm{R}}\varphi(x)g_{\frac{n}{N}}(x){\mathrm{d}}x+o_{N}(1)\\ q_{n,N}^{N}(0,\psi)&=\sum\limits_{v\in{\mathbbm{Z}}}\sqrt{N}\int_{\frac{v}{\sqrt{N}}}^{\frac{v+1}{\sqrt{N}}}\psi(y){\mathbf{P}}(S_{N-n}=v){\mathrm{d}}y=\int_{\mathbbm{R}}\varphi(y)g_{1-\frac{n}{N}}(y){\mathrm{d}}y+o_{N}(1).\end{split}$ (2.36)
@@
By (2.26), the contribution from $n\leq\varepsilon N$ and $n\geq(1-\varepsilon)N$ is bounded by a constant $C_{\varphi,\psi}$ timing $\int_{[0,\varepsilon]\cup[1-\varepsilon,1]}G_{\vartheta}(t){\mathrm{d}}t$, which converges to 0 as $\varepsilon\to 0$. The proof is completed. ∎
@@
$\sum\limits_{r=1}^{\infty}(\widehat{h})^{r}\idotsint\limits_{0<t_{1}<\cdots<t_{r}<1}g_{t_{1}}(\varphi,0)\Big(\prod\limits_{i=2}^{r}\frac{1}{\sqrt{t_{i}-t_{i-1}}}\Big)g_{1-t_{r}}(0,\psi){\mathrm{d}}t_{1},\dots\mathrm{d}t_{r},$
which is convergent by [11, Lemma B.2]. Note that as long as $h_{N}=o((\log N)^{-1})$, the asymptotics of $\sigma_{N}^{2}$ in (1.20) remains unchanged. Since $h_{N}$ only affects the mean, we can set $h_{N}\equiv 0$ without loss of generality.
2.6 Coarse graining procedure